# DiffLoRA: Generating Personalized Low-Rank Adaptation Weights with Diffusion

### v1

@@ Introduction
Figure 1: DiffLoRA uses a diffusion transformer model to predict identity-specific LoRA weights from reference images, which are then merged into SDXL for inference.
@@
Inspired by the success of low-rank adaptation (LoRA) (Hu et al. 2022) and hypernetworks (Peebles et al. 2022; Ha, Dai, and Le 2022), we propose a novel latent diffusion-based hypernetwork framework, named DiffLoRA, designed to predict the LoRA weights. Unlike previous approaches that use hypernetworks to generate weights of specific components such as multilayer perceptron (MLP) or normalization layers (Erkoç et al. 2023; Wang et al. 2024a), our framework provides a comprehensive solution for generating the LoRA weights of specific identity, guided by the reference image conditions, as illustrated in Figure 1. DiffLoRA contains contains an autoencoder for LoRA weights and a diffusion transformer model specifically tailored for LoRA. This framework synthesizes facial and overall mixed feature information of reference images utilizing a Mixture-of-Experts-inspired (MoE-inspired) gate network (Shazeer et al. 2016). During the inference phase, the diffusion model leverages these mixed features as conditional inputs to facilitate the denoising process. By reconstructing LoRA latent representations with the decoder, we can obtain the corresponding LoRA weights for specific identity adaptation in the Stable Diffusion XL (SDXL) model (Podell et al. 2023). Such a latent diffusion-based hypernetwork framework enables the prediction of a larger scale of parameters compared to other frameworks (Wang et al. 2024a). Additionally, the diffusion framework supports cross-modality generation, achieving more stable and superior results compared to other network frameworks through iterative denoising (Ruan et al. 2023; Zhang et al. 2024). By merging the LoRA weights generated by DiffLoRA into the SDXL, the model achieves personalization without additional fine-tuning, while maintaining identity consistency and preserving the original generative capabilities without introducing additional inference cost.
@@
To summarize, the main contributions of our work include:
@@
  • We propose a specialized LoRA weight autoencoder for LoRA weights compression and reconstruction. Additionally, we introduce a MOE-inspired gate network that dynamically integrates detailed facial features and general image features for enhanced identity extraction.
  • We introduce a new data pipeline for constructing the LoRA weight dataset, which facilitates DiffLoRA training and ensures the representation of a diverse range of identities.
@@
$L_{LDM}=\mathbb{E}_{t,z,\epsilon}\left[\left\|\epsilon-\epsilon_{\theta}\left(\sqrt{\bar{\alpha_{t}}}z_{0}+\sqrt{1-\bar{\alpha_{t}}}\epsilon,c,t\right)\right\|^{2}\right],$ (1)
@@
Low-rank adaptation (LoRA) significantly reduces the required parameter count through its low-rank structures. For an $m\times n$ matrix, LoRA stores two smaller matrices, resulting in a total parameter count of $(m+n)\times r$, where $r$ is the low-rank factor. This reduction in parameters lowers computational complexity during optimization, thereby enhancing training efficiency. Mathematically, the low-rank approximation constrains the parameter space, reducing the degrees of freedom from $mn$ to $(m+n)r$, where $r\ll\min(m,n)$. Additionally, the constrained distribution range of LoRA weights facilitates their compression and reconstruction. This more concentrated and stable parameter distribution improves generalization by reducing the risk of overfitting (Biderman et al. 2024). This inherent characteristic of LoRA makes it more efficient relative to full-rank MLP weights, enabling better performance with fewer computational resources.
@@
Subsequently, we train two autoencoders with identical parameters to reconstruct full-rank and low-rank features, respectively. During training, both autoencoders aim to reconstruct the original images. In the testing phase, we compare the reconstruction quality of the images by peak signal-to-noise ratio (PSNR) and structural similarity index (SSIM) metrics (Hore and Ziou 2010). The results, as shown in Figure 2, indicate that the autoencoder reconstructing the low-rank features achieves higher PSNR and SSIM values, suggesting superior reconstruction quality with the same computational cost.
@@
Figure 3: Overview of DiffLoRA. Our approach consists of two stages: LoRA weight autoencoder (LAE) and diffusion process. The LAE encodes and reconstructs LoRA weights. In the diffusion process, features extracted by the face and image encoder are combined with the Mixed Image Features (MIF) mechanism, and the resulting noisy latent is processed by a diffusion transformer to predict denoised latent. During inference, random noise and reference images are fed into DiffLoRA to generate LoRA weights.
Given reference images, DiffLoRA aims to efficiently generate accurate LoRA weights by leveraging a LoRA weight autoencoder (LAE) combined with a diffusion model. This approach circumvents the extra training cost associated with dual-branch architectures while achieving high-fidelity image generation through LoRA weight generation and weight merging. We begin by detailing the construction of the LAE to compress LoRA weights into a latent space. Next, following (Erkoç et al. 2023), we describe our diffusion transformer (DiT) model training to directly predict encoded LoRA latent representations, rather than the noise. We then introduce the Mixed Image Features (MIF) mechanism to integrate face features and image features from images, guiding the DiT model in the denoising process. Lastly, we outline the dataset construction pipeline for generating LoRA weights that accommodate multiple identities. Figure 3 provides an overview of our training and inference process.
@@
Concerning the informational features of LoRA weights, we conducted an in-depth analysis of the impact of different magnitudes of LoRA parameters on the final generated images. We sorted the parameters in LoRA by their magnitude and divided them into four proportional groups: the first group representing the large weights and the last group representing the small weights. Subsequently, we either set the large weights and small weights to zero or perturb them with proportionally scaled noise. As illustrated in Figure 4, perturbing or zeroing small weights has little effect on the generated images, while doing the same to large weights significantly degrades the identity fidelity of the images.
Figure 4: An example demonstrating the impact of perturbing or zeroing LoRA weights on generated images.
@@
$\mathcal{L}_{\text{WP}}=\frac{1}{n}\sum_{i=1}^{n}\left|x_{i}\right|\cdot\left|x_{i}-\hat{x}_{i}\right|,$ (2)
where $n$ is the number of parameters, $x_{i}$ represents the $i$-th parameter, and $\hat{x}_{i}$ represents the $i$-th reconstructed parameter. The term $\left|x_{i}\right|\cdot\left|x_{i}-\hat{x}_{i}\right|$ emphasizes the reconstruction error for larger weights. By integrating the weight-preserved loss into the training process, our approach ensures a more accurate preservation of critical information encoded in the larger weights, thereby significantly improving the quality of generated images post-reconstruction. Moreover, our LAE can compress the original LoRA weights by nearly 300 times, demonstrating an exceptional compression and reconstruction performance.
Mixed Image Features. The core concept of MIF is to leverage both facial details and general image information to better extract identity features, thereby improving the accuracy of the denoising process. Drawing inspiration from Mixture-of-Experts (MoE) structure (Shazeer et al. 2016), MIF combines image features and face features with a gate network. The overall workflow is illustrated in Figure 5.
Figure 5: MIF mechanism. Our MIF consists of a MOE-inspired gate network to combine face embedding $E_{face}$ and image embedding $E_{img}$ into mixed features for improved identity feature extraction.
We first obtain the face embedding containing facial details for identity recognition from the Face Encoder ($E_{f}$) and the image embedding including background and overall appearance from the Image Encoder ($E_{img}$) based on the reference image, and the concatenate these embeddings to obtain the mixed embedding.
By passing the mixed embedding to the Gate Network ($G$), we can obtain the Face Score ($S_{f}$) represents the importance of facial features, while the Image Score ($S_{i}$) represents the importance of general image features. And the mixed features can be calculated based on the scores. The forward process of MIF can be represented as:
$\mathbf{h}=S_{\text{img}}\odot\text{MLP}(E_{\text{img}}(I))+S_{\text{face}}\odot\text{MLP}(E_{\text{face}}(I)),$ (3)
where $S_{img},S_{face}=\text{Gate}(Z)=\text{Softmax}(f(Z))$ and $Z$ represent the mixed embedding. Here, $\odot$ denotes scalar multiplication, and the function $f$ can be a simple linear layer.
@@
here, $M_{\theta}$ denotes the DiT architecture model parameterized by $\theta$, $t$ is a time step uniformly sampled from the range $[0,T]$, and $\epsilon$ is standard Gaussian noise. The parameters $\alpha_{t}$ , $\sigma_{t}$, and $w_{t}$ are diffusion noise parameters. The term $x_{0}$ represents the latent representations, summed with the positional encoding vector, while $c$ denotes the condition integrated into the diffusion model through MIF.
@@
As illustrated in Figure 6, we develop the pipeline for generating the high-quality LoRA weight dataset. Our approach begins with collecting facial images from the FFHQ (Karras, Laine, and Aila 2019) and CelebA-HQ (Karras et al. 2018) datasets. To ensure the performance of LoRA weights, we utilize PhotoMaker (Li et al. 2024) to generate 100 distinct images for each individual at the resolution of 1024$\times$1024. Through this pipeline, we create a diverse image dataset suitable for training and deriving 100k LoRA weights.
@@
Image Filtering and Preprocessing. We employ InsightFace to calculate the facial similarity between the generated images and the original image. The top 85 images with the highest similarity scores are selected. These selected images are enriched and undergo further preprocessing, including cropping, flipping, and color transformation through various data augmentation operations.
@@
Training with DreamBooth. The preprocessed 85 images are then used to train the SDXL model with the DreamBooth method (Ruiz et al. 2023).
@@
Model Image Reward↑ DINO↑(%) CLIP-I↑(%) CLIP-T↑(%) Face Sim↑(%) FID↓ Speed↓(s)
@@
The encoder of LAE employs multiple 1D convolution kernels, sized to match the rank of the LoRA weights, to compress the weight vectors into a lower-dimensional vector. Attention layers and residual connections are integrated into the encoder to enhance its performance. The decoder uses attention layers combined with an MLP for reconstructing the compressed weights. The LAE optimization process is performed using AdamW on 4 NVIDIA 3090 GPUs for about a week, with a batch size of 4 and a learning rate of 2e-5. In the diffusion phase, we trained a 16-layer, 1454-dimension DiT model. During training, images of the same identity are randomly sampled as conditions to create mixed image features. We use AdamW with a batch size of 32 and a learning rate of 1e-4, implementing 1000 diffusion timesteps with a linear noise scheduler ranging from 0.0001 to 0.012.
@@
Following DreamBooth (Ruiz et al. 2023), we use the DINO (Caron et al. 2021) and CLIP-I (Gal et al. 2023) metrics to measure identity fidelity, and the CLIP-T metric to assess text-image consistency. However, since the CLIP-T metric may not fully capture the nuances of text-image alignment, we also compute the Image Reward score (Xu et al. 2024) to further evaluate the alignment and quality of the generated images concerning the text prompts. To further evaluate identity fidelity, we measure the face embedding similarity, referred to as Face Sim, extracted by the InsightFace model between the generated and reference images. The quality of the generated images is assessed using the FID metric, which compares the distribution of generated images with those from MS-COCO, where lower FID values indicate better image quality. We assess the inference cost by evaluating personalization speed based on whether the method involves tuning or is tuning-free. For tuning-free methods, only inference time is measured, whereas for methods requiring tuning, both training and inference times are included. Personalization speed is measured on a single NVIDIA A6000 GPU.
@@
Table 2: Ablation studies on identity fidelity metric (Face Sim, DINO), text-image consistency metrics (CLIP T).
@@
Quantitative Comparisons. In our quantitative experiments, the effectiveness and efficiency of DiffLora were evaluated using multiple metrics covering various aspects: text-image consistency, identity fidelity, generation quality, and personalization speed. As illustrated in Table 2, our method achieved the highest scores across almost all metrics. The results demonstrate that our method satisfies the requirements for generating high-quality images with a high level of identity fidelity, as evidenced by achieving the highest CLIP-I, DINO scores, and Face Sim. Additionally, Difflora exhibited the lowest inference cost, as reflected in the personalization speed, compared to other methods.
@@
As illustrated in Table 2 and Figure 8, our analysis reveals the impact of different components of the DiffLoRA method on the quality of generated images.
Impact of the Mixed Image Features (MIF). When the gate network in MIF is removed from the process, using only face features (No Image Feature) or only image features (No Face Feature) for inference, there is a significant decrease in the identity preservation of the generated images. This drastic drop suggests that MIF plays a crucial role in guiding DiffLoRA to generate accurate LoRA weights. Especially, only with image feature, the fundamental features of personalized generation are almost lost.
Impact of the Weight-Preserved Loss. We trained the LAE model without the Weight Preserved Loss (No WP Loss), which resulted in a noticeable decline in image quality. The generated faces were less similar to the input faces compared to when the WP loss was incorporated. Including the Weight Preserved Loss significantly improves the LAE’s ability to reconstruct high-quality LoRA parameters. Consequently, this enhancement boosts the Face Sim from 37.1 to 42.2.
Impact of the number of reference images. Adjusting the number of reference images (N) during inference impacts identity fidelity and facial similarity. By averaging the embeddings of multiple reference images, we find that increasing the number of images improves facial similarity metrics but reduces the CLIP I metric. This indicates a trade-off between capturing detailed features and maintaining generalization, which is crucial for generating images that are both personalized and diverse.

### v2

@@ Introduction
Figure 1: DiffLoRA employs a diffusion transformer model to predict personalized LoRA weights based on reference images, which are subsequently integrated into the SDXL model for image generation.
@@
Inspired by the success of low-rank adaptation (LoRA) (Hu et al. 2022) and hypernetworks (Peebles et al. 2022; Ha, Dai, and Le 2022), we propose a novel latent diffusion-based hypernetwork framework, named DiffLoRA, designed to predict the personalized LoRA weights. Unlike previous approaches that use hypernetworks to generate weights of specific components such as multilayer perceptron (MLP) or normalization layers (Erkoç et al. 2023; Wang et al. 2024a), our framework provides a comprehensive solution for generating the LoRA weights of specific identity, guided by the reference images, as illustrated in Figure 1. DiffLoRA contains an autoencoder for LoRA weights and a diffusion transformer model specifically tailored for LoRA latent representations. This framework synthesizes facial and overall mixed features information of reference images utilizing an Mixture-of-Experts-inspired (MoE-inspired) gate network (Shazeer et al. 2016). During the inference phase, the diffusion model leverages these mixed features as conditional inputs to facilitate the denoising process. By reconstructing the generated LoRA latent representations with the decoder, we can obtain the corresponding LoRA weights for specific identity adaptation in the Stable Diffusion XL (SDXL) (Podell et al. 2023). Such a latent diffusion-based hypernetwork framework enables the prediction of a larger scale of parameters compared to other frameworks (Wang et al. 2024a). Additionally, the diffusion framework supports cross-modality generation, achieving more stable and superior results compared to other network frameworks through iterative denoising (Ruan et al. 2023; Zhang et al. 2024; Erkoç et al. 2023). By merging the LoRA weights generated by DiffLoRA into the SDXL, the model achieves personalization without additional fine-tuning, while maintaining identity fidelity and preserving the original generative capabilities without introducing additional inference cost.
@@
The summary of our contributions is as follows:
@@
  • We propose a specialized LoRA weight autoencoder for LoRA weights compression and reconstruction. Additionally, we introduce an MoE-inspired gate network that dynamically integrates detailed facial features and general image features for enhanced identity extraction.
  • We introduce a new pipeline for constructing the LoRA weight dataset, which enhances DiffLoRA training and ensures the representation of a wide range of identities.
@@
$L_{\mathrm{LDM}}=\mathbb{E}_{t,z,\epsilon}\left[\left\|\epsilon-\epsilon_{\theta}\left(\sqrt{\bar{\alpha_{t}}}z_{0}+\sqrt{1-\bar{\alpha_{t}}}\epsilon,c,t\right)\right\|^{2}\right],$ (1)
@@
The low-rank structures of LoRA significantly reduces the required parameter count from $m\times n$ to $(m+n)\times r$ by decomposing the weight update $\Delta W\in\mathbb{R}^{m\times n}$ into two low-rank matrices $B\in\mathbb{R}^{m\times r}$ and $A\in\mathbb{R}^{r\times n}$, where $r\ll\min(m,n)$. The forward pass equation is:
$W^{\prime}(x)=W_{0}(X)+B(A(x))+b,$ (2)
where $\Delta W=BA$ is the update. As noted in (Biderman et al. 2024), the more concentrated and stable parameter distribution in LoRA contributes to improved generalization by reducing the risk of overfitting. This constrained distribution range of LoRA weights also facilitates their compression and reconstruction and makes them more efficient relative to full-rank MLP, enabling better reconstruction quality with the same computational resources.
@@
Subsequently, we train two convolutional autoencoders with identical architectures to reconstruct full-rank and low-rank features, respectively. During the training process, both autoencoders aim to reconstruct the input features. Once the autoencoder trained on low-rank features converges, we evaluate the reconstruction quality of both models using peak signal-to-noise ratio (PSNR) and structural similarity index (SSIM) metrics (Hore and Ziou 2010). The results, as shown in Figure 2, indicate that the autoencoder reconstructing the low-rank features achieves higher values in both PSNR and SSIM, demonstrating a superior reconstruction quality. In contrast, the autoencoder trained on full-rank features under the same training conditions performs significantly worse. More details of our toy experiment are provided in the supplementary materials.
@@
Figure 3: Overview of DiffLoRA. Our approach consists of two stages: LoRA weight autoencoder (LAE) and diffusion process. The LAE encodes and reconstructs LoRA weights. In the diffusion process, features extracted by the face encdoer and the image encoder are combined with the Mixed Image Features (MIF) mechanism, and the resulting noisy latent is processed by a diffusion transformer to predict denoised latent. During inference, random noise and reference images are fed into DiffLoRA to generate new LoRA weights.
Given reference images, DiffLoRA aims to efficiently generate accurate LoRA weights by leveraging a LoRA weight autoencoder (LAE) combined with a diffusion model. This approach circumvents the extra training cost associated with dual-branch architectures while achieving high-fidelity image generation through LoRA weight generation and weight merging. We begin by detailing the construction of the LAE to compress LoRA weights into the latent space. Next, following (Erkoç et al. 2023), we describe our diffusion transformer (DiT) model training to directly predict encoded LoRA latent representations, rather than the noise. We then introduce the Mixed Image Features (MIF) mechanism to integrate face features and image features from images, guiding the DiT model in the denoising process. Lastly, we outline the dataset construction pipeline for generating LoRA weights that accommodate multiple identities. Figure 3 provides an overview of our training and inference process.
@@
Concerning the informational features of LoRA weights, we conduct an in-depth analysis of the impact of different magnitudes of LoRA parameters on the final generated images. We sort the parameters in LoRA by their magnitude and divided them into four proportional groups: the first group representing the large weights and the last group representing the small weights. Subsequently, we either set the large weights and small weights to zero or perturb them with proportionally scaled noise. As illustrated in Figure 4, perturbing or setting the small weights to zero has little effect on the generated images, while doing the same to large weights significantly degrades the identity fidelity.
Figure 4: An example illustrating the impact of perturbing or setting LoRA weights to zero on generated images.
@@
$L_{\text{WP}}=\frac{1}{n}\sum_{i=1}^{n}\left|x_{i}\right|\cdot\left|x_{i}-\hat{x}_{i}\right|,$ (3)
where $n$ is the number of parameters, $x_{i}$ represents the $i$-th parameter, and $\hat{x}_{i}$ represents the $i$-th reconstructed parameter. The term $\left|x_{i}\right|\cdot\left|x_{i}-\hat{x}_{i}\right|$ emphasizes the reconstruction error for larger weights. By integrating the weight-preserved loss into the training process, our approach ensures a more accurate preservation of identity information encoded in the larger weights, thereby significantly improving the quality of generated images post-reconstruction. Moreover, our LAE can compress the original LoRA weights by nearly 300 times, demonstrating an exceptional compression and reconstruction performance.
Mixed Image Features (MIF). The core concept of MIF is to leverage both facial details and general image information to better extract identity features, thereby improving the accuracy of the denoising process. Drawing inspiration from Mixture-of-Experts (MoE) structure (Shazeer et al. 2016), MIF combines image features and face features with a gate network. The overall workflow is illustrated in Figure 5.
Figure 5: Our MIF consists of an MoE-inspired gate network to combine face embedding and image embedding into mixed features for improved identity feature extraction.
We first obtain the face embedding containing facial details for identity recognition from InsightFace encoder $E_{\text{face}}$ (Deng 2022) and the image embedding including overall appearance from CLIP image encoder $E_{\text{img}}$ (Radford et al. 2021) based on the reference image I, and then concatenate these embeddings to obtain the mixed embedding containing more comprehensive identity information.
By passing the mixed embedding $M$ to the Gate Network $G$, we can obtain the Face Score $S_{\text{face}}$ represents the importance of facial features, while the Image Score $S_{\text{img}}$ represents the importance of general image features. And the mixed features $F$ can be calculated based on the scores. The forward process of MIF can be represented as:
$F=S_{\text{img}}\odot MLP(E_{\text{img}}(I))+S_{\text{face}}\odot MLP(E_{\text{face}}(I)),$ (4)
where $\odot$ denotes scalar multiplication, $S_{\text{img}}$ and $S_{\text{face}}$ are the outputs of the Gate Network $G(M)$. Specifically, $G(M)=\text{Softmax}(f(M))$. The function $f$ can be a linear layer.
@@
here, $M_{\theta}$ denotes the DiT architecture model parameterized by $\theta$, $t$ is a time step uniformly sampled from the range $[0,T]$, and $\epsilon$ is standard Gaussian noise. The parameters $\alpha_{t}$ , $\sigma_{t}$, and $w_{t}$ are diffusion noise parameters. The term $x_{0}$ represents the original latent representations, while $c$ denotes the condition integrated into the diffusion model through MIF.
@@
As illustrated in Figure 6, we develop the pipeline for generating the high-quality LoRA weight dataset. Our approach begins with collecting facial images from the FFHQ (Karras, Laine, and Aila 2019) and CelebA-HQ (Karras et al. 2018) datasets. To ensure the performance of LoRA weights, we utilize PhotoMaker (Li et al. 2024) to generate 100 distinct images for each individual at the resolution of 1024$\times$1024. Through this pipeline, we create a diverse image dataset suitable for training and deriving 100k LoRA checkpoints.
@@
Image Filtering and Preprocessing. We employ InsightFace (Deng 2022) to calculate the facial similarity between the generated images and the original image. The top 85 images with the highest similarity scores are selected. These selected images are enriched and undergo further preprocessing, including cropping, flipping, and color transformation through various data augmentation operations.
@@
Training with DreamBooth. The 85 preprocessed images are then used to train the SDXL model following the method in (Ruiz et al. 2023) and to apply LoRA as described in (Yeh et al. 2023). This process fine-tunes the SDXL and derives the LoRA weights for high-fidelity image generation.
@@
Method Image Reward↑ DINO↑(%) CLIP-I↑(%) CLIP-T↑(%) Face Sim↑(%) FID↓ Speed↓(s)
@@
The encoder of LAE employs multiple 1D convolutional layers to compress the flatten LoRA weight vectors into a lower-dimensional vector. Attention layers and residual connections are integrated into the encoder to enhance its performance. The decoder uses attention layers combined with an MLP for reconstructing the compressed LoRA weights. The LAE optimization process is performed using AdamW on 4 NVIDIA A6000 GPUs for about a week, with a batch size of 4 and a learning rate of 2e-5. In the diffusion phase, we trained a 16-layer, 1454-dimension DiT model. During training, images of the same identity are randomly sampled as conditions to create mixed image features. We use AdamW with a batch size of 32 and a learning rate of 1e-4, implementing 1000 diffusion timesteps with a linear noise scheduler ranging from 0.0001 to 0.012. Training runs for approximately 1400 epochs until convergence, taking about 2 days on 4 NVIDIA A6000 GPUs. Inference is conducted using 100 steps of the DDIM sampler.
@@
Following DreamBooth (Ruiz et al. 2023), we use the DINO (Caron et al. 2021) and CLIP-I (Gal et al. 2023) metrics to measure identity fidelity, and the CLIP-T metric to assess text-image consistency. However, since the CLIP-T metric may not fully capture the nuances of text-image alignment, we also compute the Image Reward score (Xu et al. 2024) for a more comprehensive evaluation of the generated images relative to the text prompts. Additionally, to further evaluate identity fidelity, we measure the face embedding similarity, referred to as Face Sim, extracted by the InsightFace between the generated and reference images. The quality of the generated images is assessed using the FID metric (Heusel et al. 2017), which compares the distribution of generated images with those from MS-COCO (Lin et al. 2014), where lower FID values indicate better image quality. We assess the inference cost by evaluating personalization speed based on whether the method involves tuning or is tuning-free. For tuning-free methods, only inference time is measured, whereas for methods requiring tuning, both training and inference times are included. Personalization speed is measured on a single NVIDIA A6000 GPU.
@@
Table 2: Ablation studies on identity fidelity metric (Face Sim, DINO, CLIP-I).
@@
Quantitative Comparisons. In our quantitative experiments, the effectiveness and efficiency of DiffLoRA were evaluated using multiple metrics covering various aspects: text-image consistency, identity fidelity, generation quality, and personalization speed. As illustrated in Table 2, our method achieved the highest scores across almost all metrics. The results demonstrate that our method satisfies the requirements for generating high-quality images with a high level of identity fidelity, as evidenced by achieving the highest CLIP-I, DINO and Face Sim scores. Additionally, DiffLoRA exhibited the lowest inference cost, as reflected in the personalization speed, compared to other methods.
@@
As illustrated in Table 2 and Figure 8, our analysis reveals the impact of different components of the DiffLoRA method on the quality of generated images with the specific identity. More details are listed in the supplementary materials.
Impact of the Mixed Image Features (MIF). When the gate network in MIF is removed from the process, using only face features (No Image Feature) or only image features (No Face Feature) for inference, there is a significant decrease in the identity preservation of the generated images. This drastic drop suggests that MIF plays a crucial role in guiding DiffLoRA to generate accurate LoRA weights. Especially, only with image features, the identity features of personalized generation are almost lost.
Impact of the Weight-Preserved Loss. We train the LAE without the weight-preserved loss (No WP Loss), which results in a noticeable decline in image quality. The generated identity are less similar to the input identity without weight-preserved loss. Incorporating the weight-preserved loss significantly improves the LAE’s ability to reconstruct more accurate LoRA parameters. Consequently, this enhancement boosts the Face Sim from 37.31 to 41.84.
Impact of the Number of Reference Images. Adjusting the number of reference images (N) utilized during the inference process exhibits varying impacts on identity fidelity. We achieve inference with multiple reference images by averaging their embeddings. Observations indicate that as the number of reference images increases, the model captures more comprehensive identity information, leading to an enhancement in identity fidelity metric.